### v1

@@ 3.4 The Algorithm
Let $T=\Theta(\log^{2}n)$ and $c=\Theta(\frac{1}{\phi\log n})$. The algorithm for Theorem 3.1 is presented in Algorithm 1. The algorithm runs for at most $T$ rounds and stops when $\mu(R_{t})>\frac{\mu(V)\cdot c\cdot\phi}{70}=\Omega(\frac{\mu(V)}{\log n})$. In each round $t$, we implicitly update $F_{t}$ (see Section 3.2). In order to keep the running time near linear, we do the flow computation using the fair cut result of [7]. These routines may also return a cut $S_{t}\subseteq A_{t}$ with $\Phi_{G[A]}(S_{t},A_{t}\setminus S_{t})\leq\frac{7}{c}$, in which case we “move” $S_{t}$ from $A_{t+1}$ to $R_{t+1}$. After $T$ rounds, $F_{T}$ certifies that the remaining part of $A_{T}$ is a near $(\Omega(\phi),\mu)$-expander.
@@
Our key technical Lemmas 4.8 and 4.9 carefully establish the decrease in potential achieved by the cut player, taking into account the matchings and the deleted cuts. We conclude in Corollary 4.11 that after $O(\log^{2}n)$ iterations the potential falls below $1/n$. We then prove, using our variation on Cheeger’s inequality (see Lemma 4.12), that having a small potential means that $A_{t}$ is a near $(\Omega(1),\mu)$-expander in $F_{t}$.
@@ 4.2 $A_{T}$ is a near expander in $F_{T}$
$\mathop{\mathbb{E}}\left[\frac{1}{3}\sum_{\{i,k\}\in M_{t}}w_{ik}\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right\|_{2}^{2}+\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)\left\|\frac{W_{t}(j)}{\sqrt{\mu(j)}}\right\|_{2}^{2}\right]\geq\frac{1}{3000\alpha\log n}\psi(t)-\frac{3}{n^{\alpha/16}}$
@@
and $z=\max(\mathcal{B})$. Let $K\coloneqq\langle\mathbbm{1},\mu\rangle=poly(m)$. We get that
@@
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{\geq}}\frac{n}{3\alpha\log{n}}\sum_{i\in A_{t}^{l}\setminus S_{t}}m_{i}(u_{i}-\eta)^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in S_{t}}\mu(j)u_{j}^{2}-2K\cdot z$
$\geq\frac{n}{3\alpha\log{n}}\sum_{i\in A_{t}^{l}\setminus S_{t}}m_{i}(u_{i}-\eta)^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in A_{t}^{l}\cap S_{t}}\mu(j)u_{j}^{2}-2K\cdot z$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu4\mkern-1.5mu)}}{{\geq}}\frac{n}{27\alpha\log{n}}\sum_{i\in A_{t}^{l}\setminus S_{t}}m_{i}u_{i}^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in A_{t}^{l}\cap S_{t}}\mu(j)u_{j}^{2}-2K\cdot z$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu5\mkern-1.5mu)}}{{\geq}}\frac{n}{27\alpha\log{n}}\sum_{i\in A_{t}^{l}}m_{i}u_{i}^{2}-2K\cdot z\geq\frac{n}{3000\alpha\log{n}}\sum_{i\in A_{t}}\mu(i)u_{i}^{2}-2K\cdot z\ .$
@@
We note that $z=O(poly(n))$. Indeed, for every $i\in A_{t}\cap\mathcal{T}$
$u_{i}=\left\langle\frac{W_{t}(i)}{\sqrt{\mu(i)}},r\right\rangle\leq\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}\right\|_{2}=\frac{\left\|W_{t}(i)\right\|_{2}}{\sqrt{\mu(i)}}\leq\frac{1}{\sqrt{\mu(i)}}=O(poly(n)),$
@@
By Lemma C.3, $z$ is non-zero with probability at most $\frac{1}{n^{\alpha/8}}$, so for large enough $\alpha$, we get $2K\mathop{\mathbb{E}}[z]\leq\frac{2Kpoly(n)}{n^{\alpha/8}}\leq\frac{3}{n^{\alpha/16}}$, [^10] The second Inequality explains why Assumption 2.1 requires $\min_{v\in\mathcal{T}}\mu(v)\leq poly(n)$. completing the proof. ∎
@@
For each round $t$, $\mathop{\mathbb{E}}[\psi(t+1)]\leq\left(1-\frac{1}{3000\alpha\log n}\right)\psi(t)+\frac{3}{n^{\alpha/16}}$, where the expectation is over the unit vector $r\in\mathop{\mathbb{R}}^{n}$.
@@
In order to prove near-expansion we need to lower bound $|E_{F_{T}}(S,\bar{S})|$. We do so by upper bounding $|E_{F_{T}}(S,S)|=\mathbbm{1}_{S}^{\prime}F_{T}\mathbbm{1}_{S}$. Because $F_{t}$ is $\mathcal{T}$-blocked we get that $\mathbbm{1}_{S}^{\prime}F_{T}\mathbbm{1}_{S}=\mathbbm{1}_{S}^{\prime}(I_{T}F_{T}I_{T})\mathbbm{1}_{S}$. Observe the following relation between $X_{T}$ and $I_{T}F_{T}I_{T}$:
@@
So $\Phi_{G}(S,\bar{S})=\frac{|E(S,\bar{S})|}{\mu(S)}\geq\frac{1}{5}$, and this is true for all cuts $S\subseteq A$ with $\frac{\mu(S)}{\mu(A_{t})}\leq\frac{1}{2}$.
@@
Recall that $S_{t}$ denotes the cut returned by Lemma 3.6 (applied on $G[A_{t}]$) at iteration $t$, so that $A_{t+1}=A_{t}\setminus S_{t}$.
Observe first that in any round $t$, we have $\Phi^{\mu}_{G}(A_{t},R_{t})\leq\frac{7}{c}=O(\phi\log n)$. This is because $R_{t}=\bigcup_{0\leq t^{\prime}<t}{S_{t^{\prime}}}$ and by Lemma 3.6, for each $t^{\prime}$, $\Phi^{\mu}_{G[A_{t^{\prime}}]}(S_{t^{\prime}},\bar{S_{t^{\prime}}})\leq\frac{7}{c}=O(\phi\log n)$.
Assume Algorithm 1 terminates because $\mu(R_{t})>\frac{\mu(V)\cdot c\cdot\phi}{70}=\Omega(\frac{\mu(V)}{\log n})$. We also have, by Lemma 3.6, that $\mu(A_{t})=\Omega(\mu(V))=\Omega(\frac{\mu(V)}{\log n})$. Then $(A_{t},R_{t})$ is a balanced cut where $\Phi_{G}(A_{t},R_{t})=O(\phi\log n)$. We end in Case (2) of Theorem 3.1.
Otherwise, Algorithm 1 terminates at round $T$ and we apply Corollary 4.13. If $R=\emptyset$, then we obtain the first case of Theorem 3.1 because the whole vertex set $V$ is, with high probability, a near $(\Omega(\phi),\mu)$-expander, which means that $G$ is an $(\Omega(\phi),\mu)$-expander. Otherwise, we write $c=\frac{c_{1}}{\phi\log n}$ for some constant $c_{1}$, and $c_{0}\mathrel{\mathop{:}}=\frac{7}{c_{1}}$. We have $\Phi_{G}(A_{T},R_{T})\leq\frac{7}{c}=\frac{7}{c_{1}}\phi\log n=c_{0}\phi\log n$. Additionally, $\mu(R_{t})\leq\frac{\mu(V)\cdot c\cdot\phi}{70}=\frac{\mu(V)\cdot c_{1}}{70\log n}=\frac{\mu(V)}{10c_{0}\log n}$, and, with high probability, $A_{T}$ is a near $(\Omega(\phi),\mu)$-expander in $G$, which means we obtain the third case of Theorem 3.1.
@@
    Constructing $M_{t}$ in time $O(m\log n)$ (using dynamic trees [11]).
@@ 5 Expander Decomposition
The following lemma, called the trimming step, is the last key component for the expander decomposition algorithm. This is a slight generalization of Theorem 8.2 of [7]. For completeness, we include its proof in Appendix A for completeness.
@@
Given a graph $G=(V,E)$, $A\subseteq V$, a parameter $\phi>0$, and a weight function $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$ over the vertices, such that $A$ is a near $(\phi,\mu)$-expander in $G$ and $|E(A,\bar{A})|\leq\frac{\phi\cdot\mu(A)}{9}$ The trimming step finds with high probability $A^{\prime}\subseteq A$ in time $\tilde{O}(m)$ such that $\Phi^{\mu}(G[A^{\prime}])\geq\phi/6$. Moreover, $\mu(A^{\prime})\geq\mu(A)-\frac{4|E(A,\bar{A})|}{\phi}$, and $|E(A^{\prime},\bar{A^{\prime}})|\leq 2|E(A,\bar{A})|$.
@@
Given a graph $G=(V,E)$ of $m$ edges, a parameter $\phi>0$, and a weight function $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$ over the vertices, there exists a randomized algorithm which takes $\tilde{O}(m)$ time and must end in one of the following three cases:
@@
    We certify that $G$ has $\mu$-expansion $\Phi^{\mu}(G)=\Omega(\phi)$, with high probability.
@@
    We find a cut $(A,\bar{A})$ in $G$ of $\mu$-expansion $\Phi^{\mu}_{G}(A,\bar{A})=O(\phi\log n)$, and $\mu(A),\mu(\bar{A})$ are both $\Omega\left(\frac{\mu(V)}{\log n}\right)$, i.e, we find a relatively balanced low $\mu$-expansion cut.
@@
    We find a cut $(A,\bar{A})$ with $0<\mu(\bar{A})\leq\frac{\mu(V)}{\log n}$, $\Phi^{\mu}_{G}(A,\bar{A})=O(\phi\log n)$, and with high probability, $G[A]$ is $\Omega(\phi)$-expander with respect to $\mu$.
@@
We apply Theorem 3.1. Cases (1) and (2) translate directly to Cases (1) and (2) of Theorem 5.2. If Theorem 3.1 terminates with Case (3), we use Theorem 5.1. We have $\mu(\bar{A})\leq\frac{\mu(V)}{10c_{0}\log n}$, and $\Phi_{G}^{\mu}(A,\bar{A})\leq c_{0}\phi\log n$. This means that $|E(A,\bar{A})|\leq\frac{\mu(V)\cdot\phi}{10}\leq\frac{\mu(A)\cdot\phi}{9}$. The trimming step takes $\tilde{O}\left(m\right)$ time, and we return the cut $(A^{\prime},\bar{A^{\prime}})$. It satisfies that $G[A^{\prime}]$ is a $\Omega(\phi)$ expander with respect to $\mu$. Additionally, $\mu(\bar{A^{\prime}})\leq\mu(\bar{A})+\mu(A)-\mu(A^{\prime})\leq\mu(\bar{A})+\frac{4|E(A,\bar{A})|}{\phi}\leq\frac{\mu(V)}{10c_{0}\log n}+\frac{2\mu(V)}{5}\leq\frac{\mu(V)}{2}$. This means that $\bar{A^{\prime}}$ is still the smaller side of the cut, with respect to $\mu$. By Theorem 5.1, we have that $|E(A^{\prime},\bar{A^{\prime}})|\leq 2|E(A,\bar{A})|$, so $\Phi_{G}^{\mu}(A^{\prime},\bar{A^{\prime}})=\frac{|E(A^{\prime},\bar{A^{\prime}})|}{\mu(\bar{A^{\prime}})}\leq 2\frac{|E(A,\bar{A})|}{\mu(\bar{A})}=2\Phi_{G}^{\mu}(A,\bar{A})=O(\phi\log n)$. If $\mu(\bar{A^{\prime}})\leq\frac{\mu(V)}{\log n}$, we can terminate with Case (3) of the theorem ($\mu(\bar{A^{\prime}})\geq\mu(\bar{A})>0$). Otherwise, $\mu(A^{\prime})$, $\mu(\bar{A^{\prime}})$ are both $\Omega\left(\frac{\mu(V)}{\log n}\right)$, so we can terminate with case (2) of the theorem. ∎
We now show how to get an expander decomposition from Theorem 5.2. The following is the expander decomposition procedure of [12].
Algorithm 2 Expander Decomposition [12]
@@
2:  Call Cut-Matching($G$, $\phi$) $\triangleright$ See Theorem 5.2
3:  if we certify that $\Phi^{\mu}_{G}\geq\phi$ then return $G$
4:  else if we find a relatively balanced cut $(A,R)$ then
5:   return Decomp($G[A]$, $\phi$, $\mu$) $\cup$ Decomp($G[R]$, $\phi$, $\mu$)
6:  else we find a very unbalanced cut $(A,R)$ $\triangleright$ This cut was trimmed - AA is an expander
7:   return Decomp($G[R]$, $\phi$, $\mu$) $\cup A$
@@
To bound the running time, note that if we get Case (2) of Theorem 5.2 then both sides of the cut have total $\mu(\cdot)$ measure of at most $\left(1-\Omega\left(\frac{1}{\log n}\right)\right)\cdot\mu(V)$. If we get Case (3), then $\mu(A)=\Omega(\mu(V))$, so the total $\mu(\cdot)$ is reduced by a constant factor (i.e., $\mu(\bar{A})\leq\left(1-\Omega(1)\right)\cdot\mu(V)$). In any case, the total $\mu(\cdot)$ measure of the largest component drops by a factor of at least $1-\Omega\left(\frac{1}{\log n}\right)$ across each level of the recursion, so (as $\mu(V)=O(poly(n))$, by Assumption 2.1) the recursion depth is $O(\log^{2}n)$. [^12] To be precise, the recursion depth is $O(\log n\log\mu^{\star})$, where $\mu^{\star}=\frac{\max(\mu(\cdot))}{\min(\mu(\cdot))}$.
@@
To bound the number of edges between expander clusters, observe that in both Case (2) and Case (3), we always cut a component along a cut of $\mu$-expansion $O(\phi\log n)$. Thus, we can charge the edges on the cut to the $\mu(\cdot)$ of the vertices on the smaller side of the cut (with respect to $\mu$), so each vertex $v$ is charged $O(\phi\mu(v)\log n)$. A vertex can be on the smaller side of a cut at most $O(\log n)$ times, so we can charge each vertex $v$ at most [^13] To be precise, we can charge each vertex at most $O(\phi\mu(v)\log n\log\mu^{\star})$. $O(\phi\mu(v)\log^{2}n)$ to pay for all the edges between the final clusters. This bounds the total number of edges between the expanders to be at most $O(\phi\mu(V)\log^{2}n)$. ∎

### v2

@@ 3.4 The Algorithm
Let $T=\Theta(\log^{2}n)$ and $c=\Theta(\frac{1}{\phi\log n})$. The algorithm for Theorem 3.1 is presented in Algorithm 1. The algorithm runs for at most $T$ rounds and stops when $\mu(R_{t})>\frac{\mu(V)\cdot c\cdot\phi}{70}=\Omega(\frac{\mu(V)}{\log n})$. In each round $t$, we implicitly update $F_{t}$ (see Section 3.2). In order to keep the running time near linear, we compute the flow using the fair cut result of [LNPSsoda13]. This routine may also return a cut $S_{t}\subseteq A_{t}$ which satisfies $\Phi_{G[A]}(S_{t},A_{t}\setminus S_{t})\leq\frac{7}{c}$ (with high probability), in which case we “move” $S_{t}$ from $A_{t+1}$ to $R_{t+1}$. After $T$ rounds, $F_{T}$ certifies that the remaining part of $A_{T}$ is a near $(\Omega(\phi),\mu)$-expander with high probability.
@@
Our key technical Lemmas 4.8 and 4.9 carefully establish the decrease in potential achieved by the cut player, taking into account the matchings and the deleted cuts. We conclude in Corollary 4.11 that after $O(\log^{2}n)$ iterations the potential falls below $1/n$ with high probability. We then prove, using our variation on Cheeger’s inequality (see Lemma 4.12), that having a small potential means that $A_{t}$ is a near $(\Omega(1),\mu)$-expander in $F_{t}$.
@@ 4.2 $A_{T}$ is a near expander in $F_{T}$
$\mathop{\mathbb{E}}\left[\frac{1}{3}\sum_{\{i,k\}\in M_{t}}w_{ik}\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right\|_{2}^{2}+\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)\left\|\frac{W_{t}(j)}{\sqrt{\mu(j)}}\right\|_{2}^{2}\right]\geq\frac{1}{3000\alpha\log n}\psi(t)-\frac{1}{n^{\alpha/16}}$
@@
and $z=\max(\mathcal{B})$. Let $K\coloneqq\langle\mathbbm{1},\mu\rangle=\operatorname{\mathbf{poly}}(n)$. We get that
@@
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{\geq}}\frac{n}{3\alpha\log{n}}\sum_{i\in A_{t}^{l}\setminus S_{t}}m_{t}(i)(u_{i}-\eta)^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in S_{t}}\mu(j)u_{j}^{2}-2K\cdot z$
$\geq\frac{n}{3\alpha\log{n}}\sum_{i\in A_{t}^{l}\setminus S_{t}}m_{t}(i)(u_{i}-\eta)^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in A_{t}^{l}\cap S_{t}}\mu(j)u_{j}^{2}-2K\cdot z$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu4\mkern-1.5mu)}}{{\geq}}\frac{n}{27\alpha\log{n}}\sum_{i\in A_{t}^{l}\setminus S_{t}}m_{t}(i)u_{i}^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in A_{t}^{l}\cap S_{t}}\mu(j)u_{j}^{2}-2K\cdot z$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu5\mkern-1.5mu)}}{{\geq}}\frac{n}{27\alpha\log{n}}\sum_{i\in A_{t}^{l}}m_{t}(i)u_{i}^{2}-2K\cdot z\geq\frac{n}{3000\alpha\log{n}}\sum_{i\in A_{t}}\mu(i)u_{i}^{2}-2K\cdot z\ .$
@@
We note that $z=O(\operatorname{\mathbf{poly}}(n))$. Indeed, for every $i\in A_{t}\cap\mathcal{T}$
$u_{i}=\left\langle\frac{W_{t}(i)}{\sqrt{\mu(i)}},r\right\rangle\leq\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}\right\|_{2}=\frac{\left\|W_{t}(i)\right\|_{2}}{\sqrt{\mu(i)}}\leq\frac{1}{\sqrt{\mu(i)}}=O(\operatorname{\mathbf{poly}}(n)),$
@@
By Lemma C.3, $z$ is non-zero with probability at most $\frac{1}{n^{\alpha/8}}$, so for large enough $\alpha$, we get $2K\mathop{\mathbb{E}}[z]\leq\frac{2K\operatorname{\mathbf{poly}}(n)}{n^{\alpha/8}}\leq\frac{1}{n^{\alpha/16}}$, [^12] The second Inequality explains why Assumption 2.1 requires $\max_{v\in\mathcal{T}}\mu(v)\leq\operatorname{\mathbf{poly}}(n)$. completing the proof. ∎
@@
For each round $t$, $\mathop{\mathbb{E}}[\psi(t+1)]\leq\left(1-\frac{1}{3000\alpha\log n}\right)\psi(t)+\frac{1}{n^{\alpha/16}}$, where the expectation is over the unit vector $r\in\mathop{\mathbb{R}}^{n}$.
@@
In order to prove near-expansion we need to lower bound $|E_{F_{T}}(S,\overline{S})|$. We do so by upper bounding $|E_{F_{T}}(S,S)|=\mathbbm{1}_{S}^{\prime}F_{T}\mathbbm{1}_{S}$. Because $F_{t}$ is $\mathcal{T}$-blocked we get that $\mathbbm{1}_{S}^{\prime}F_{T}\mathbbm{1}_{S}=\mathbbm{1}_{S}^{\prime}(I_{T}F_{T}I_{T})\mathbbm{1}_{S}$. Observe the following relation between $X_{T}$ and $I_{T}F_{T}I_{T}$:
@@
So $\Phi_{G}(S,\overline{S})=\frac{|E(S,\overline{S})|}{\mu(S)}\geq\frac{1}{5}$, and this is true for all cuts $S\subseteq A$ with $\frac{\mu(S)}{\mu(A_{t})}\leq\frac{1}{2}$.
@@
Recall that $S_{t}$ denotes the cut returned by Lemma 3.6 (applied on $G[A_{t}]$) at iteration $t$, so that $A_{t+1}=A_{t}\setminus S_{t}$. We write $c_{1}\mathrel{\mathop{:}}=c\phi\log n=O(1)$, and let $c_{0}\mathrel{\mathop{:}}=\frac{7}{c_{1}}$.
Observe first that in any round $t$, we have $\Phi^{\mu}_{G}(A_{t},R_{t})\leq\frac{7}{c}=O(\phi\log n)$. Indeed, since $R_{t}=\bigcup_{0\leq t^{\prime}<t}{S_{t^{\prime}}}$, Lemma 3.6 implies that, with high probability, for each $t^{\prime}$, $\Phi^{\mu}_{G[A_{t^{\prime}}]}(S_{t^{\prime}},\bar{S_{t^{\prime}}})\leq\frac{7}{c}=O(\phi\log n)$.
Assume Algorithm 1 terminates because $\mu(R_{t})>\frac{\mu(V)\cdot c\cdot\phi}{70}=\Omega(\frac{\mu(V)}{\log n})$. Hence, $\mu(R_{t-1})\leq\frac{\mu(V)\cdot c\cdot\phi}{70}=\frac{\mu(V)\cdot c_{1}}{70\log n}$. By Lemma 3.6, we also have $\mu(S_{t})\leq\frac{2}{3}\mu(A_{t})\leq\frac{2}{3}\mu(V)$ with high probability. Combining these inequalities gives $\mu(R_{t})=\mu(R_{t-1})+\mu(S_{t})\leq\frac{3}{4}\mu(V)$ [^14] For sufficiently large $n$, we have $\log n\geq c_{1}$.. In particular, $\mu(A_{t})=\mu(V)-\mu(R_{t})=\Omega(\mu(V))=\Omega(\frac{\mu(V)}{\log n})$. Therefore $(A_{t},R_{t})$ is a balanced cut where $\Phi_{G}(A_{t},R_{t})=O(\phi\log n)$. We end in Case (2) of Theorem 3.1.
Otherwise, Algorithm 1 terminates at round $T$ and we apply Corollary 4.13. If $R=\emptyset$, then we obtain the first case of Theorem 3.1 because the whole vertex set $V$ is, with high probability, a near $(\Omega(\phi),\mu)$-expander, which means that $G$ is an $(\Omega(\phi),\mu)$-expander. Otherwise, we have $\Phi_{G}(A_{T},R_{T})\leq\frac{7}{c}=\frac{7}{c_{1}}\phi\log n=c_{0}\phi\log n$. Additionally, $\mu(R_{t})\leq\frac{\mu(V)\cdot c\cdot\phi}{70}=\frac{\mu(V)\cdot c_{1}}{70\log n}=\frac{\mu(V)}{10c_{0}\log n}$, and, with high probability, $A_{T}$ is a near $(\Omega(\phi),\mu)$-expander in $G$, which means we obtain the third case of Theorem 3.1.
@@
    Constructing $M_{t}$ in time $O(m\log n)$ (using dynamic trees [ST83]).
@@ 5 Expander Decomposition
The following lemma, called “the trimming step”, is the last key component for the expander decomposition algorithm. This is a slight generalization of [LNPSsoda13, Theorem 8.2]. For completeness, we include its proof in Appendix A.
@@
Given a graph $G=(V,E)$, a set $A\subseteq V$, a parameter $\phi>0$, and a vertex measure $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$, such that $A$ is a near $(\phi,\mu)$-expander in $G$ and $|E(A,\overline{A})|\leq\frac{\phi\cdot\mu(A)}{9}$, there exists a randomized algorithm (the trimming step) which finds in time $\tilde{O}(m)$ a set $A^{\prime}\subseteq A$ such that with high probability:
  • $\Phi^{\mu}(G[A^{\prime}])\geq\phi/6$,
  • $\mu(A^{\prime})\geq\mu(A)-\frac{4|E(A,\overline{A})|}{\phi}$, and
  • $|E(A^{\prime},\overline{A^{\prime}})|\leq 2|E(A,\overline{A})|$.
@@
Given a graph $G=(V,E)$ with $n$ vertices and $m$ edges, a parameter $\phi>0$, and a vertex measure $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$ satisfying Assumption 2.1, there exists a randomized algorithm which takes $\tilde{O}(m)$ time and must end in one of the following three cases:
@@
    We certify that $G$ has $\mu$-expansion $\Phi^{\mu}(G)=\Omega(\phi)$.
@@
    We find a cut $(A,\overline{A})$ in $G$ of $\mu$-expansion $\Phi^{\mu}_{G}(A,\overline{A})=O(\phi\log n)$, and $\mu(A),\mu(\overline{A})$ are both $\Omega\!\left(\frac{\mu(V)}{\log n}\right)$, i.e, we find a relatively balanced low $\mu$-expansion cut.
@@
    We find a cut $(A,\overline{A})$ with $0<\mu(\bar{A})\leq\frac{\mu(V)}{2}$, $\Phi^{\mu}_{G}(A,\overline{A})=O(\phi\log n)$, and $G[A]$ is $\Omega(\phi)$-expander with respect to $\mu$.
The correctness of the algorithm holds with high probability.
@@
We apply Theorem 3.1. Cases (1) and (2) translate directly to Cases (1) and (2) of Theorem 5.2. If Theorem 3.1 terminates with Case (3), we use Theorem 5.1. We have $\mu(\overline{A})\leq\frac{\mu(V)}{10c_{0}\log n}$, and $\Phi_{G}^{\mu}(A,\overline{A})\leq c_{0}\phi\log n$. This means that for large enough $n$, $|E(A,\overline{A})|\leq\frac{\mu(V)\cdot\phi}{10}\leq\frac{\mu(A)\cdot\phi}{9}$. The trimming step takes $\tilde{O}\left(m\right)$ time, and we return the cut $(A^{\prime},\overline{A^{\prime}})$ as Case (3). Indeed, $G[A^{\prime}]$ is a $\Omega(\phi)$ expander with respect to $\mu$. Additionally, for large enough n, $\mu(\overline{A^{\prime}})=\mu(\overline{A})+\mu(A)-\mu(A^{\prime})\leq\mu(\overline{A})+\frac{4|E(A,\overline{A})|}{\phi}\leq\frac{\mu(V)}{10c_{0}\log n}+\frac{2\mu(V)}{5}\leq\frac{\mu(V)}{2}$. This means that $\overline{A^{\prime}}$ is still the smaller side of the cut, with respect to $\mu$. By Theorem 5.1, we have that $|E(A^{\prime},\overline{A^{\prime}})|\leq 2|E(A,\overline{A})|$, so $\Phi_{G}^{\mu}(A^{\prime},\overline{A^{\prime}})=\frac{|E(A^{\prime},\overline{A^{\prime}})|}{\mu(\overline{A^{\prime}})}\leq 2\frac{|E(A,\overline{A})|}{\mu(\overline{A})}=2\Phi_{G}^{\mu}(A,\overline{A})=O(\phi\log n)$. ∎
Note that case $(3)$ may return a balanced cut. We get the following stronger version of Theorem 5.2 by classifying case $(3)$ as case $(2)$ whenever $\mu(\overline{A})>\frac{\mu(V)}{\log n}$.
Corollary 5.3.
Given a graph $G=(V,E)$ with $n$ vertices and $m$ edges, a parameter $\phi>0$, and a vertex measure $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$ satisfying Assumption 2.1, there exists a randomized algorithm which takes $\tilde{O}(m)$ time and must end in one of the following three cases:
  1. 1.
    We certify that $G$ has $\mu$-expansion $\Phi^{\mu}(G)=\Omega(\phi)$.
  2. 2.
    We find a cut $(A,\overline{A})$ in $G$ of $\mu$-expansion $\Phi^{\mu}_{G}(A,\overline{A})=O(\phi\log n)$, and $\mu(A),\mu(\overline{A})$ are both $\Omega\!\left(\frac{\mu(V)}{\log n}\right)$, i.e, we find a relatively balanced low $\mu$-expansion cut.
  3. 3.
    We find a cut $(A,\overline{A})$ with $0<\mu(\bar{A})\leq\frac{\mu(V)}{\log n}$, $\Phi^{\mu}_{G}(A,\overline{A})=O(\phi\log n)$, and $G[A]$ is $\Omega(\phi)$-expander with respect to $\mu$.
The correctness of the algorithm holds with high probability.
Proof.
In case Theorem 5.2 terminated with Case (3) and the cut was relatively balanced (specifically, $\frac{\mu(V)}{\log n}<\mu(\overline{A})\leq\frac{\mu(V)}{2}$), we instead terminate with Case (2). All conditions of Case (2) are satisfied. ∎
We now show how to get an expander decomposition from Theorem 5.2. The following is based on the expander decomposition procedure of [saranurak2019expander, agassy2022expander].
Algorithm 2 Expander Decomposition [saranurak2019expander, agassy2022expander]
@@
2:  Call Cut-Matching($G$, $\phi$) $\triangleright$ See Corollary 5.3.
3:  if Case (1): we certify that $\Phi^{\mu}_{G}\geq\phi$ then return $\{V\}$.
4:  else if Case (2): we find a relatively balanced cut $(A,\overline{A})$ then
5:   return Decomp($G[A]$, $\phi$, $\mu$) $\cup$ Decomp($G[\overline{A}]$, $\phi$, $\mu$).
6:  else Case (3): we find a very unbalanced cut $(A,\overline{A})$ $\triangleright$ $G[A]$ is an expander.
7:   return Decomp($G[\overline{A}]$, $\phi$, $\mu$) $\cup\{A\}$.
@@
To bound the running time, note that if we get Case (2) of Corollary 5.3 then both sides of the cut have total $\mu(\cdot)$ measure of at most $\left(1-\Omega\!\left(\frac{1}{\log n}\right)\!\right)\!\cdot\!\mu(V)$. If we get Case (3), then $\mu(A)=\Omega(\mu(V))$, so the total $\mu(\cdot)$ is reduced by a constant factor (i.e., $\mu(\overline{A})\leq\left(1-\Omega(1)\right)\cdot\mu(V)$). In any case, the total $\mu(\cdot)$ measure of the largest component drops by a factor of at least $1-\Omega\!\left(\frac{1}{\log n}\right)$ across each level of the recursion, so (as $\mu(V)=O(\operatorname{\mathbf{poly}}(n))$, by Assumption 2.1) the recursion depth is $O(\log^{2}n)$.
@@
To bound the number of edges between expander clusters, observe that in both Case (2) and Case (3), we always cut a component along a cut of $\mu$-expansion $O(\phi\log n)$. Thus, we can charge the edges on the cut to the $\mu(\cdot)$ of the vertices on the smaller side of the cut (with respect to $\mu$), so each vertex $v$ is charged $O(\phi\mu(v)\log n)$. A vertex can be on the smaller side of a cut at most $O(\log n)$ times, so we can charge each vertex $v$ at most $O(\phi\mu(v)\log^{2}n)$ to pay for all the edges between the final clusters. This bounds the total number of edges between the expanders to be at most $O(\phi\mu(V)\log^{2}n)$. ∎